OpenStudy (tennistar):

A company's monthly profit, P, from a product is given by P = -x2 + 105x - 1050, where x is the price of the product in dollars. What is the lowest price of the product, in dollars, that gives a monthly profit of $1,550? (Do not put a $ in your answer.)

OpenStudy (jackellyn):

1500= -x^2 +105x -1050 -x^2 +105x -1050-1500 -x^2 +105x -2550 do the quadratic formula a=-1 b=105 c=-2550 can you?

OpenStudy (tennistar):

no

OpenStudy (tennistar):

@rebeccaskell94

OpenStudy (tennistar):

plz help me understand

OpenStudy (accessdenied):

Essentially, we want to find the smallest value of \(x\) that makes our monthly profit \(P = 1550\). To do this, we may let \(P = 1550\). Then, we plug in the equivalent expression of \(P\) in terms of \(x\) given by the problem (\(P = -x^2 + 105x - 1050\)). So our equation is: \(-x^2 + 105x -1050 = 1550\) Does that part make sense?

OpenStudy (tennistar):

yeah, I got the answer, with a little help thanks

OpenStudy (accessdenied):

Ah, okay. You're welcome. What answer did you get, just checking?